OpenStudy (anonymous):

i want to verify my answer please.. .. find y' if y=sin^-1(e^-2x)

OpenStudy (anonymous):

my answer is \[y'=\left(\begin{matrix}1 \\ \sqrt{1+x ^{2}}\end{matrix}\right)\left( e ^{-2x} \right)+\sin^{-1} \left( -2xe ^{-2x} \right)\]

OpenStudy (nikvist):

do you think \[y=\arcsin{e^{-2x}}\]

OpenStudy (anonymous):

i think is the same

OpenStudy (anonymous):

ohhhh

OpenStudy (anonymous):

upsss

OpenStudy (anonymous):

its is wrong

OpenStudy (nikvist):

\[y'=\frac{1}{\sqrt{1-\left(e^{-2x}\right)^2}}\cdot\left(e^{-2x}\right)'\]

OpenStudy (anonymous):

ohhhh :)

OpenStudy (anonymous):

the chain rule for that will be -2xe^-2x